How much force is needed for a 2700-pound car to accelerate at 3.5 feet-per-second-squared? Neglect all other forces such as aerodynamic drag. Enter the answer in pounds
Answer #1

- 1 found = 2700 pound = 0.4536 kg 1225 lg | 1 f/2 = 0.3 m/42 3.5 frize = 1007 m/s2 F=ma = 1225*1.07 = 1310.75 N 11 Newton =